An air filled with 75 microfarad has a voltage of 12V. An insulator with dielectric constant of 1.00095 is inserted between the metal plates. Calculate the capacitance. What is the new voltage across the capacitor? Calculate the original charge on the capacitor. How much charge is currently on the capacitor


1
Expert's answer
2022-04-27T13:34:34-0400

The capacitance

"C=\kappa C_0\\
=1.00095*75\:\rm \mu F=75.07\:\mu F"

New voltage

"V=V_0\frac{C_0}{C}\\
=\rm 12\: V*1/1.00095=11.99\: V"

Original charge on the capacitor

"q_0=C_0V_0\\
=\rm 75*10^{-6}\: F*12\: V=9.0*10^{-4}\: C"